OpenStudy (twizttiez):

Dosha is creating a new dessert that has two layers shaped like cones. The inner cone is frozen ice cream and has a diameter of 12 cm and a height of 6 cm. The outer layer is a thin wafer shell, like an upside-down ice cream cone, with a height of 15 cm and the same diameter as the inner layer. Dosha will inject a cream filling into the space. What is the volume of the cream filling? Use 3.14 to approximate pi and express your final answer in hundredths.

OpenStudy (twizttiez):

Can you help @Reaper534

OpenStudy (twizttiez):

@confluxepic

OpenStudy (twizttiez):

Sorry I forgot to tag you oops XD

OpenStudy (twizttiez):

@TheSmartOne ?

OpenStudy (twizttiez):

Can you help?

TheSmartOne (thesmartone):

Find the volume of the bigger cone. Find the volume of the smaller cone. Answer = Volume of bigger cone - Voulme of smaller cone

OpenStudy (twizttiez):

How do I do that? Like I don't know the formula :/

TheSmartOne (thesmartone):

Volume of a cone = \(\Large\sf \frac{1}{3}\times \pi \times r^2\times h\)

OpenStudy (twizttiez):

What would the 1 third be for?

OpenStudy (twizttiez):

?

TheSmartOne (thesmartone):

Well the formula has 1/3 in it.. Kinda hard to explain .-. here: http://math.stackexchange.com/questions/623/why-is-the-volume-of-a-cone-one-third-of-the-volume-of-a-cylinder

OpenStudy (twizttiez):

Okay so where would I place my numbers?

TheSmartOne (thesmartone):

So first it tells us that: h = 15 d= 12 Hint: \(\sf\Large\frac{d}{2}=r\) So what is r?

OpenStudy (twizttiez):

6 right?

OpenStudy (twizttiez):

So I have 1/3*pi*6^2*15 right?

TheSmartOne (thesmartone):

correct!

TheSmartOne (thesmartone):

os is lagging so bad, sorry about the dalays >.<